# Downsizing Diffusion Models for Cardinality Estimation

### v1

@@ 5. The Selectivity Estimator
5.1. Predictor-Corrector Integration: Why GMM
While the score estimator and density estimator are effective in calculating the probability density function’s value at any given point, we still need to integrate this function over the entire queried region in order to estimate the selectivity of ranged queries. ADC does this by utilizing the importance sampling Monte Carlo formula
$\int_{V}p(x)dx=\int_{V}q(x)dx\cdot\mathrm{E}_{x\sim q(x|x\in V)}\frac{p(x)}{q(x)}$
that Naru (Yang et al., 2019) and DQM-D(Hasan et al., 2020)’s selectivity estimator are both based on, which, when the expectation on the right is approximated using i.i.d. sampling, becomes an unbiased estimator with variance proportionate to
@@ 5.1. Predictor-Corrector Integration: Why GMM
To enhance speed and precision, the selectivity estimator of ADC must construct a predictor that can be rapidly integrated, rapidly sampled from, and resembles the corrector $p_{\epsilon,\theta}$ as much as possible. However, the predictor constructed by Naru (via progressive sampling) and DQM-D (via VEGAS (Lepage, 1978)) require multiple phases of sampling. This is undesirable for us, considering that evaluating the density function at one point already requires a not-so-easy numerical integration process.
@@
In terms of similarity to $p$, another advantage of GMM is that the underlying distribution $p_{\epsilon}$, which $p_{\epsilon,\theta}$ seeks to approximate, always have a KDE-type expression
@@
as in the database case, $p_{0}$ is simply the average of delta functions placed at sample points. Given their unique shape and smoothness features, ”Gaussian mixtures work best at matching local features of Gaussian mixtures” is a very reasonable assumption.
@@
using the EM algorithm (Dempster, 1977), an elegant GMM-exclusive optimizer that trains GMMs by iteratively alternating between estimating the posterior responsibilities of mixture components (E-step) and updating the model parameters to maximize the expected complete-data log-likelihood (M-step). Compared to one-size-fit-all optimizers like Stochastic Gradient Descent (SGD), EM is very robust and efficient, as it requires no hyperparameters, guarantees monotone increase of likelihood, and always converges to a local maximum value.
ADC further attempts to escape from local maximums, something the classical EM algorithm is incapable of, by resampling kernels with extremely low weights. That is, after every 10 EM iterations, ADC deletes all Gaussian kernels with weights below $1/500N$, and replace them with new kernels of weight $1/500N$, covariance $k^{2}I_{d}$, and mean values resampled from $\{x_{j}\}_{j=1}^{M}$ via a weighted sampler that sets the weight of a data point $x_{i}$ inversely proportionate to the current value of $q(x_{i})$. If training time is sufficient, it is also sometimes preferable to train multiple GMMs at once, and then select the model that achieves the highest likelihood.
5.2. Neutralizing Early Stopping Error
We briefly discuss two ways of using our GMM predictor-corrector scheme to cancel out the error caused by early stopping, shown by experiments to be crucial for ADC’s optimal performance.
Our first method is based on the following observations: First, $q$, when trained to maximize the function (8), is in fact trying to approximate not $p_{\epsilon}$, but $p_{0}=\frac{1}{M}\sum_{i=1}^{M}\delta(x_{i})$ itself. Second, perturbing GMMs using a diffusion process is extremely easy: setting $q_{0}(x)$ to be the $q(x)$ defined in (7), we would simply have
@@ 5.2. Neutralizing Early Stopping Error
Our second method is based on the observation that both GMMs and diffusion models work by ”bleeding” portions of density from the typical set onto neighboring regions. When one or more pairs of attributes $(A_{i},A_{j})$ exhibit near-functional dependency, hence the typical set’s projection onto $(A_{i},A_{j})$ becomes a 1-dimensional curve, such a practice significantly distorts the local probability distribution, leading to large errors for extreme-case queries.
ADC mitigates this problem by using Bayesnet (Chow and Liu, 1968) to handle functionally dependent attribute pairs. In case where knowing the value of an attribute $A_{i}$ can help narrow down the range of another attribute $A_{j}$ to $k$ (chosen to be 0.05 in the current model) times the original, ADC, instead of training a GMM and a diffusion model on all attributes, would instead work by training:
@@
  • a Bayesnet histogram modeling the conditional probability density $p(A_{j}|A_{i})$ as a function piecewise constant with respect to $A_{i}$ and $A_{j}$.
@@
$\sum_{m=1}^{N}w_{m}\left(\prod_{k\neq{i,j}}\int_{a_{k}}^{b_{k}}\varphi_{H_{m,kk}}(x-y_{m,k})dx\cdot g(m,a_{j},b_{j})\right),$
with $g(m,a_{j},b_{j})$ defined as
$\int_{a_{i}}^{b_{i}}\varphi_{H_{m,ii}}(x-y_{m,i})p(y_{j}\in[a_{j},b_{j}]|x_{i})dx_{i},$
then corrects its prediction using the weighed average of $\{\frac{p_{\epsilon}(z_{k})}{q_{\epsilon}(z_{k})}\}_{k=1}^{n}$, with $\{z_{k}\}_{k=1}^{n}$ i.i.d. samples from $q_{\epsilon}$ and the weight of each sample point set proportionate to
$\{p(x_{j}\in[a_{j},b_{j}]|z_{k,i})\}_{k=1}^{n}.$
Because $p$, constructed using histograms, is constant with respect to $x_{i}$ in each of the intervals $\{[x_{i,p},x_{i,p+1}]\}_{p=0}^{P-1}$, calculation of the above expressions can be greatly sped up by calculating and caching, for the dimension $i$, each slice $x_{i,p},x_{i,p+1}]$, and each Gaussian kernel $m$,
$\int_{x_{i,p}}^{x_{i,p+1}}\varphi_{H_{m,ii}}(z_{m,i}-x_{i})dx_{i}$
once upon initializing the model; as well as calculating
$\{p(y_{j}\in[p_{j},q_{j}]|x_{i}\in[x_{i,p},x_{i,p+1}])\}_{p=0}^{P-1}$
@@
For our benchmark datasets, the above Bayesnet implementation worked very well on power, successfully cutting down the max Q-error of queries from 100 to 28, as the range of the 0th attribute can be narrowed down to an average of 4.5% its original by knowing the 4th attribute’s value.
@@
8:  Sample $\{x_{i}\}_{i=1}^{n}\sim q_{\epsilon}(x|x\in V)$
9:  $\{y_{i}\}_{i=1}^{n}\leftarrow\{\frac{p_{s_{\theta},\epsilon}(x_{i})}{q_{\epsilon}(x_{i})}\}_{i=1}^{n}$
10:  $Sel\leftarrow Q\cdot\mathrm{mean\{y_{i}\}_{i=1}^{n}}$
@@
5.3. Introducing ADC+
Considering that the advantages of ADC+ are more easily verified experimentally instead of theoretically, more information regarding our motives behind creating ADC+, as well as the performance gains resulting from it, shall be provided in 6.3.
@@ 6.1.1. Datasets
We test the performance of ADC on two real world and one synthetic dataset, including:
Forest(Asuncion et al., 2007): Forest coverage type dataset with 581012 rows, used as a benchmark by multiple influential works on cardinality estimation (Wang et al., 2021) (Heimel et al., 2015) (Kipf et al., 2018) (Dutt et al., 2019). Following the practice of (Wang et al., 2021), we keep the first 10 continuous attributes for evaluation.
Power(Asuncion et al., 2007): Dataset with 2.1 million rows, recording power consumption information of a French household at a 1 minute sampling rate, used alongside forest in all four essays above. Following the practice of (Wang et al., 2021), we keep the 7 continuous measurement attributes excluding date and time, and convert ”?” to -1 when dealing with the 1.25% of rows consisting entirely of missing values.
Modulo: A synthetic dataset designed according to the philosophy that lead to our creation of ADC: treating all attributes as a single, unified entity is at times necessary for producing decent estimates. The dataset consists of 4 million rows and five attributes AA,$B$,$C$,$D$, and $E$, where $A$, $B$, and $D$ are independent integers taking the values 0 to 1999 with equal probability;
@@
and $\epsilon_{1}$, $\epsilon_{2}$ are independent integers taking the values 0 to 199 with equal probability. A property of modulo is that all five attributes are pairwise independent from one another, making it potentially challenging for estimators that work by discerning pairwise correlations.
@@ 6.1.2. Workloads
For each dataset, we create 100,000 queries for training query-driven methods, 10,000 queries for validation, and 10,000 queries for testing using the program provided by (Wang et al., 2021). That is, for a $d$-dimensional dataset, we design queries asking for all points lying within a hyperrectagle, where:
@@ 6.1.3. Comparative Techniques
We briefly introduce versions of our model and the models we use for comparison, which include all learned estimators tested in (Wang et al., 2021). For the sake of giving each model a similar size budget while preventing poor results due to bad parameter choices, comparative techniques are trained using the same hyperparameter settings the author of (Wang et al., 2021) chose for the dataset forest, putting the parameter size of each model at around 0.5MB per dataset.
@@
Pred: The performance of ADC’s GMM predictor as a standalone estimator, which we report to to quantify accuracy improvements gained by the diffusion model corrector, i.e. the score estimator and density estimator.
MSCN (Kipf et al., 2018): a query-driven regression model which processes the feature vector of queries using a multi-set convolutional network.
Lw-Tree and LW-NN (Dutt et al., 2019): two lightweight query-driven regression models that process the feature vector of queries using either a gradient boosted tree (LW-Tree) or a neural network (LW-NN), and enhance their precision with the help of CE (correlation based) features extracted from simple heuristic estimators.
DeepDB (Hilprecht et al., 2020): a data-driven joint distribution model that reconstructs data distributions using sum-product networks.
Naru (Yang et al., 2019): a data-driven joint distribution model that reconstructs data distributions by training an autoregressive model and answers ranged queries using sequential sampling, shown by (Wang et al., 2021) to be the most robust and accurate learned model on most test datasets.
@@ 6.1.4. Experiment Device and Accuracy Metric
We run all our experiments on a dual-socket Intel Xeon Gold 6140 system with 36 physical cores (72 hardware threads) and a base clock frequency of 2.30 GHz. We choose Q-error, defined as
@@
to be our accuracy metric, with zero values for $Card_{real}$ and $Card_{est}$ both set to 1 to prevent division by zero. We report $\max$, 99th percentile, 95th percentile, median, and geometric mean Q-error as our evaluation benchmarks in accordance with (Wang et al., 2021).
We note that the performance of LW-Tree/LW-NN on forest and MSCN, Naru, DeepDB on power are not as decent in our experiments as they were in (Wang et al., 2021)’s, possibly due to different Python library versions our smaller model size budget for the power dataset [^4] In a personal communication, professor Xiaoying Wang, the original author of (Wang et al., 2021), confirmed that such deviations are plausible and likely attributable to stochastic differences arising from software library versions. They also said that in their original experiments, their GM Q-error results were 1.32 for LW-Tree and 1.35 for LW-NN. While we record in Table 1 the results from our experiments for methodological consistency, we also encourage readers to look up the original figures in (Wang et al., 2021) for a more comprehensive comparison.
6.2. Comparative Experiment Evaluation
@@
(a) Q-error of Different Estimators on Real-world Datasets (”Pred” is ADC’s GMM Predictor)
Models Forest Power
GM 50th 95th 99th max GM 50th 95th 99th max
ADC+ 1.30 1.10 3.00 7.00 511 1.05 1.01 1.26 2.02 267
ADC 1.32 1.15 3.00 7.00 511 1.09 1.06 1.26 2.00 288
Pred 1.36 1.11 4.00 9.00 2037 1.05 1.01 1.27 2.22 267
MSCN 1.72 1.17 8.00 19.5 138 1.26 1.02 5.00 19.9 392
DeepDB 1.35 1.06 4.25 13.0 3408 1.09 1.00 1.53 3.88 1870
LW-NN 1.51 1.22 4.37 14.33 8306 1.17 1.06 1.89 4.25 $2.2\cdot 10^{4}$
LW-Tree 1.64 1.33 5.50 13.90 3176 1.13 1.02 1.75 4.33 $2.6\cdot 10^{4}$
Naru 1.27 1.06 3.21 8.00 452 1.05 1.01 1.19 2.00 3174
@@
(c) Latency (L,ms) and Storage Cost (S,MB)
Models Forest Power Modulo
L S L S L S
ADC+ 24 0.34 9.4 0.26 7.3 0.22
MSCN 0.49 0.61 0.36 0.54 0.46 0.50
DeepDB 14 0.56 4.8 0.54 4.3 0.67
LW-NN 0.26 0.43 0.33 0.43 0.33 0.42
LW-Tree 0.12 0.65555The open-source code of (Wang et al., 2021) did not report LW-Tree’s parameter size like it did with the other four models; thus, the parameter size of LW-Tree is estimated under the assumption that LW-Tree’s full storage cost/parameter size ratio is roughly identical to that of LW-NN. 0.12 0.60 0.12 0.59
Naru 51 0.65 25.8 0.38 34.1 0.43

### v2

@@ 5. The Selectivity Estimator
5.1. Predictor Choice: Why GMM
While the score estimator and density estimator are effective in calculating the probability density function’s value at any given point, we still need to integrate this function over the entire queried region in order to estimate the selectivity of ranged queries. ADC does this by utilizing the importance sampling Monte Carlo formula (1) that Naru (refsix) and DQM-D(refseven)’s selectivity estimator are both based on, which, when $\mathrm{E}_{q(x|x\in V)}\frac{p(x)}{q(x)}$ is approximated using i.i.d. sampling, becomes an unbiased estimator with variance proportionate to
@@ 5.1. Predictor Choice: Why GMM
To enhance speed and precision, the selectivity estimator of ADC must construct a predictor that can be rapidly integrated, rapidly sampled from, and resembles the corrector $p_{\epsilon,\theta}$ as much as possible. However, the predictors constructed by Naru (via progressive sampling) and DQM-D (via VEGAS (refeleven)) require multiple phases of sampling. This is undesirable for us, considering that evaluating the density function at one point already requires a not-so-easy numerical integration process.
@@
In terms of similarity to $p$, another advantage of GMM is that the underlying distribution $p_{\epsilon}$, which $p_{\epsilon,\theta}$ seeks to approximate, always has a KDE(reffourteen)-type expression
@@
as in the database case, $p_{0}$ is simply the average of delta functions placed at sample points. Given their unique shape and smoothness features, we can reasonably assume that Gaussian mixtures work best at matching local features of Gaussian mixtures.
@@
using the EM algorithm (refthirtyseven), an elegant GMM-exclusive optimizer that trains GMMs by iteratively alternating between estimating the posterior responsibilities of mixture components (E-step) and updating the model parameters to maximize the expected complete-data log-likelihood (M-step). Compared to one-size-fits-all optimizers like Stochastic Gradient Descent (SGD), EM is very robust and efficient, as it requires no hyperparameters, guarantees a monotone increase of likelihood, and always converges to a local maximum value.
An important detail of our GMM implementation is based on the following observations: first, $q$, when trained to maximize the function (11), is in fact trying to approximate $p_{0}=\frac{1}{M}\sum_{i=1}^{M}\delta(x_{i})$ itself rather than the perturbed $p_{\epsilon}$. Second, perturbing GMMs using a diffusion process is extremely easy: setting $q_{0}(x)$ to be the $q(x)$ defined in (10), we would simply have
@@
While a somewhat bold approach, numerical experiments have proven such a design crucial for ADC’s optimal performance.
5.2. A GMM-Bayesnet Hybrid
A common flaw shared by ADC’s predictor and corrector is that GMMs and diffusion models both work by ”bleeding” portions of density from the typical set onto neighboring regions. When one or more pairs of attributes $(A_{i},A_{j})$ exhibit near-functional dependency, hence the typical set’s projection onto $(A_{i},A_{j})$ becomes a 1-dimensional curve, such a practice significantly distorts the local probability distribution, leading to large errors for extreme-case queries.
Thus, ADC further introduces Bayesnet (refsixteen) to handle functionally dependent attribute pairs. In the case where knowing the value of an attribute $A_{i}$ can help narrow down the range of another attribute $A_{j}$ to $c$ (chosen to be 0.05 in the current model) times the original, ADC, instead of training a GMM and a diffusion model on all attributes, would instead work by training:
@@ 5.2. A GMM-Bayesnet Hybrid
  • a Bayesnet histogram modeling the conditional probability density $\tilde{p}(A_{j}|A_{i})$ as a function piecewise constant with respect to $A_{i}$ and $A_{j}$.
@@
$\sum_{m=1}^{N}w_{m}\left(\prod_{k\neq{i,j}}\int_{a_{k}}^{b_{k}}\varphi_{H_{m,kk}}(x_{k}-y_{m,k})dx_{k}\cdot g_{i}(m,a_{j},b_{j})\right),$
with $g_{i}(m,a_{j},b_{j})$ defined as
$\int_{a_{i}}^{b_{i}}\varphi_{H_{m,ii}}(x_{i}-y_{m,i})\tilde{p}(y_{j}\in[a_{j},b_{j}]|x_{i})dx_{i},$
then corrects its prediction using the weighed average of $\{\frac{p_{\epsilon}(z_{r})}{q_{\epsilon}(z_{r})}\}_{r=1}^{n}$, with $\{z_{r}\}_{r=1}^{n}$ i.i.d. samples from $q_{\epsilon}$ and the weight of each sample point set proportionate to
$\{\tilde{p}(x_{j}\in[a_{j},b_{j}]|z_{r,i})\}_{r=1}^{n}.$
Because $\tilde{p}$, constructed using histograms, is constant with respect to $x_{i}$ in each of the intervals $\{[x_{i,s},x_{i,s+1}]\}_{s=0}^{S-1}$, with $S$ the number of partition intervals for the $i$th attribute’s range, the above formulas can be evaluated much faster by calculating and caching, for the dimension $i$, each slice $x_{i,s},x_{i,s+1}]$, and each Gaussian kernel $m$,
$\int_{x_{i,s}}^{x_{i,s+1}}\varphi_{H_{m,ii}}(z_{m,i}-x_{i})dx_{i}$
once upon initializing the model, as well as calculating
$\{\tilde{p}(y_{j}\in[a_{j},b_{j}]|x_{i}\in[x_{i,s},x_{i,s+1}])\}_{s=0}^{S-1}$
@@
For our benchmark datasets, the above Bayesnet implementation worked very well on power and higgs, successfully cutting down the max Q-error of queries by more than five times, as the range of the 0th attribute can be narrowed down to an average of 4.5% of its original by knowing the 4th attribute’s value.
5.3. Introducing ADC+
As the advantages of ADC+ are more easily verified experimentally instead of theoretically, more information regarding our motives behind creating ADC+, as well as the performance gains resulting from it, shall be provided in subsection 6.3.
@@ 5.3. Introducing ADC+
8:  Sample $\{x_{r}\}_{r=1}^{n}\sim q_{\epsilon}(x|x\in V)$
9:  $\{y_{r}\}_{r=1}^{n}\leftarrow\{\frac{p_{s_{\theta},\epsilon}(x_{r})}{q_{\epsilon}(x_{r})}\}_{r=1}^{n}$
10:  $Sel\leftarrow Q\cdot\mathrm{mean}\{y_{r}\}_{r=1}^{n}$
@@ 6.1.1. Datasets
We test the performance of ADC on three real-world and one synthetic dataset, including:
Forest(reftwentythree): Forest coverage type dataset with 581012 rows, used as a benchmark by multiple influential works on cardinality estimation (refeighteen) (reffourteen) (refseventeen) (reffive). Following the practice of (reffive), we keep the first 10 continuous attributes for evaluation.
Power(reftwentythree): Dataset with 2.1 million rows, recording power consumption information of a French household at a 1-minute sampling rate, used alongside forest in all four essays above. Following the practice of (reffive), we keep the 7 continuous measurement attributes excluding date and time, and convert ”?” to -1 when dealing with the 1.25% of rows consisting entirely of missing values.
Higgs(reftwentythree): A dataset with 11 million rows and 28 continuous attributes, recording the kinematic properties of microscopic particles measured by detectors in an accelerator, used as a cardinality estimation benchmark by (refthirtyfive) and used by (refthirtysix) to train their summarization model in identifying common frequency and correlation patterns. Following the practice of (refthirtyfive), we use the 7 high-level features to construct our test dataset  [^4] We round all values in higgs to three-digit decimals because Naru, due to the encoding strategy chosen by (reffive), can take weeks for training and minutes for inference when given the original data. We note that ADC (which lacks an encoder) and most other models can train and infer just as smoothly on the original dataset..
Modulo: A synthetic dataset designed according to the philosophy that leads to our creation of ADC: treating all attributes as a single, unified entity is at times necessary for producing decent estimates. The dataset consists of 4 million rows and five attributes AA,$B$,$C$,$D$, and $E$, where $A$, $B$, and $D$ are independent integers taking the values 0 to 1999 with equal probability;
@@
and $\epsilon_{1}$, $\epsilon_{2}$ are independent integers taking the values 0 to 199 with equal probability. A property of modulo is that all five attributes are pairwise independent from one another, making it challenging for estimators that work by discerning pairwise correlations.
We note that modulo is artificially and specifically designed to highlight a correlation pattern that conventional estimators fail to capture, and the results on this dataset should best be interpreted as ”We found a rare but solid loophole that no existing model, except for ours, managed to address”, rather than ”ADC can outperform other models in real-world scenarios as much as it did here”.
@@ 6.1.2. Workloads
For each dataset, we create 100,000 queries for training query-driven methods, 10,000 queries for validation, and 10,000 queries for testing using the program provided by (reffive). That is, for a $d$-dimensional dataset, we design queries asking for all points lying within a hyperrectangle, where:
@@ 6.1.3. Comparative Techniques
We briefly introduce versions of our model and the models we use for comparison, which include all learned estimators tested in (reffive).
For the sake of giving each model a similar size budget while preventing poor results due to bad parameter choices, comparative techniques are trained using the same hyperparameter settings the author of (reffive) chose for the dataset forest, putting the parameter size of each model at around 0.5MB per dataset except Naru on higgs (due to its expensive encoding scheme).
@@
  • ADC-: The performance of ADC’s selectivity estimator as a standalone model, which we report to quantify accuracy improvements gained by the diffusion model corrector, i.e., the score estimator and density estimator.
  • MSCN (refseventeen): a query-driven regression model which processes the feature vector of queries using a multi-set convolutional network.
  • Lw-Tree and LW-NN (refeighteen): two lightweight query-driven regression models that process the feature vector of queries using either a gradient boosted tree (LW-Tree) or a neural network (LW-NN), and enhance their precision with the help of CE (correlation based) features extracted from simple heuristic estimators.
  • DeepDB (refthirtythree): a data-driven joint distribution model that reconstructs data distributions using sum-product networks.
  • Naru (refsix): a data-driven joint distribution model that reconstructs data distributions by training an autoregressive model and answers ranged queries using sequential sampling, shown by (reffive) to be the most robust and accurate learned model on most test datasets.
@@ 6.1.4. Experiment Device and Accuracy Metric
We run all of our experiments on a server equipped with dual-socket Intel Xeon Gold 6140 systems, each Intel Xeon Gold 6140 having 36 physical cores (72 hardware threads) and a base clock frequency of 2.30 GHz. We choose Q-error, defined as
@@
to be our accuracy metric, with zero values for $Card_{real}$ and $Card_{est}$ both set to 1 to prevent division by zero. We report $\max$, 99th percentile, 95th percentile, median, and geometric mean Q-error as our evaluation benchmarks in accordance with (reffive).
We note that the performance of LW-Tree/LW-NN on forest and MSCN, Naru, DeepDB on power are not as good in our experiments as they were in (reffive)’s, possibly due to different Python library versions and our smaller size budget for the ”power” dataset  [^5] In a personal communication, professor Xiaoying Wang, the original author of (reffive), confirmed that such deviations are plausible and likely attributable to stochastic differences arising from software library versions. They also said that in their original experiments, their GM Q-error results were 1.32 for LW-Tree and 1.35 for LW-NN. While we record in Table 2 the results from our experiments for methodological consistency, we also encourage readers to look up the original figures in (reffive) for a more comprehensive comparison.
6.2. Experimental Results
@@
(a) Q-Error of Different Estimators on Real-world Datasets
Models Forest Power Higgs
GM 50th 95th 99th max GM 50th 95th 99th max GM 50th 95th 99th max
MSCN 1.72 1.17 8.00 19.5 138 1.26 1.02 5.00 19.9 392 1.20 1.01 2.99 14.8 199
DeepDB 1.35 1.06 4.25 13.0 3408 1.09 1.00 1.53 3.88 1870 1.06 1.00 1.28 2.77 866
LW-NN 1.51 1.22 4.37 14.33 8306 1.17 1.06 1.89 4.25 $2.2\cdot 10^{4}$ 1.10 1.02 1.50 3.57 $2.6\cdot 10^{4}$
LW-Tree 1.64 1.33 5.50 13.90 3176 1.13 1.02 1.75 4.33 $2.6\cdot 10^{4}$ 1.09 1.01 1.52 4.17 928
Naru 1.27 1.06 3.21 8.00 452 1.05 1.01 1.19 2.00 3174 1.04 1.00 1.11 2.06 258
ADC- (ours) 1.35 1.11 4.00 8.00 1978 1.05 1.00 1.23 2.24 190 1.03 1.00 1.13 1.86 143
ADC (ours) 1.32 1.16 3.00 6.00 426 1.07 1.03 1.27 2.00 128 1.05 1.01 1.14 2.00 78
ADC+ (ours) 1.28 1.10 3.00 6.00 426 1.04 1.00 1.22 2.00 128 1.03 1.00 1.13 1.96 78
@@
ADC- 1.11 1.01 1.74 3.12 70
ADC 1.12 1.07 1.45 2.06 40
ADC+ 1.08 1.01 1.44 2.06 40
(c) Latency (L,ms), Training Time (T,min), and Storage Cost (S,MB)
Forest Power Higgs Modulo
L T S L T S L T S L T S
0.49 20 0.61 0.36 20 0.54 0.37 20 0.54 0.46 20 0.50
14 12 0.56 4.8 5.7 0.54 4.4 6.9 1.5 4.3 6.2 0.67
0.26 102 0.43 0.33 103 0.43 0.27 102 0.43 0.33 104 0.42
0.12 0.18 0.65666The open-source code of (reffive) did not report LW-Tree’s parameter size like it did with the other four models; thus, the parameter size of LW-Tree is estimated under the assumption that LW-Tree’s full storage cost/parameter size ratio is roughly identical to that of LW-NN. 0.12 0.15 0.60 0.12 0.18 0.57 0.12 0.16 0.59
51 49 0.65 26 116 0.38 155 1770 1.6 34 156 0.43
3.0 600 0.11 4.9 600 0.11 4.4 600 0.11 2.7 600 0.07
34 600 0.44 23 600 0.33 23 600 0.33 16 600 0.22
26 600 0.44 7.9 600 0.33 4.8 600 0.33 7.2 600 0.22